Coffee drinkers. Among 100 coffee drinkers, 36 said they preferred their coffee black and the rest did not prefer their coffee black. Find the ratio of those who prefer black coffee to those who prefer nonblack coffee.
Answer by checkley77(12844)

36/(100-36)=36/64=9/16 ans.